# Projecting Molecules into Synthesizable Chemical Spaces

### v1

@@ 3.5 Experiment Setup
We use the SynNet reaction template set (Gao et al., 2021) curated by the authors from two publicly available template sets (Hartenfeller et al., 2012; Button et al., 2019). The template set contains 91 reaction templates encoded in SMARTS string format (Daylight Chemical Information Systems, Inc., 2023), of which 13 are uni-molecular reactions and 78 are bi-molecular reactions.
@@
We use the building blocks in the Enamine US Stock catalog retrieved in October 2023 (Enamine, 2023). For building blocks that include more than one molecule (e.g., salts, hydrates), we keep the largest molecule and drop the remaining ones. Building blocks that fail the RDKit sanitization check or do not match any reaction template are removed. Duplicate building blocks are also removed. The pre-processing procedure leads to the final building block set containing 211,220 molecules.
@@
To evaluate the similarity between the input molecule and the output molecule, we use the Tanimoto similarity score on three different fingerprints: (1) Morgan fingerprint of length 4096 and radius 2 (Morgan, 1965), (2) Morgan fingerprint of Murcko scaffold, and (3) Gobbi pharmacophore fingerprint (Gobbi & Poppinger, 1998). The three similarity scores indicate chemical similarities in three different aspects: overall structure, scaffold structure, and pharmacophore property, respectively, and are all normalized to $[0,1]$.
@@ 4.1 Generalization in Bottom-Up Synthesis Planning
We first validate our model’s ability in bottom-up synthesis planning, particularly in proposing synthetic pathways for novel molecular structures. To assess this capability with unseen molecules, we employed the following testing approach. First, we clustered building blocks into 128 groups using K-means algorithm based on their fingerprints, reserving one cluster exclusively for testing while using the remaining 127 for training. This approach generated a test set of 1,000 molecules from the reserved building block to evaluate our model’s generalization. During testing, we allow the model to access the reserved building blocks. In addition, we include another challenging benchmark: 1,000 molecules from the ChEMBL database (Release 33)(Gaulton et al., 2012), reported to be predominantly “unreachable” in previous work (Gao et al., 2021). This not only tested our model against complex, unseen molecular structures but also set a strenuous standard for assessing its generalizability.
We evaluate the model using the following metrics: (1) success rate: the percentage of valid postfix notations; (2) reconstruction rate: the percentage of proposed synthetic pathways that lead to products identical to input molecules; (3) average similarities. The similarity score is set to 0 for failed cases. We compare our model with the baseline model SynNet (Gao et al., 2021).
@@ 4.2 Projecting Molecules Generated by Structure-Based Drug Design Models
Generative models for structure-based drug design (SBDD) build molecules to fit a protein pocket structure (Luo et al., 2021). SBDD models often produce chemically invalid structures, because they operate in the 3D space with a few chemical constraints (Harris et al., 2023). As a remedy, we apply our model to molecules proposed by the SBDD models to find valid and synthesizable analogs.
We use Pocket2Mol (Peng et al., 2022) to generate molecules for targets in the LIT-PCBA dataset (Tran-Nguyen et al., 2020). The LIT-PCBA dataset contains 15 protein targets, each of which has at least one X-ray structure that can be used to run Pocket2Mol to generate molecules. For each target, we select 300 generated molecules based on the sum of QED and SA scores. Then, we apply our model to generate 5 analogs for each molecule and select best one according to Vina scores (Eberhardt et al., 2021). The estimated binding energy in kcal/mol shows slight improvement for the majority of targets (Table 3), suggesting that the top-1 synthesizable analogs of each molecule could maintain binding. Figure 4 presents two examples in which the synthesizable analogs improve predicted binding. Figure 6 and 7 in the appendix present more examples.
@@ 4.3 Projecting Molecules Generated by Goal-Directed Generative Models
Goal-directed generative models are pivotal in molecular design, yet their practicality is often hampered by the synthesizability of the produced molecules. This issue was highlighted by Gao & Coley, who, utilizing three algorithms on the GuacaMol benchmark suite’s multi-property objectives (Brown et al., 2019), revealed that around 70% of the molecules generated were considered unsynthesizable by the ASKCOS retrosynthesis analysis tool (Coley et al., 2019b).
@@
In conclusion, the projection generally leads to synthesizable analogs that are sub-optimal in terms of the oracle property scores. These sub-optimal analogs share some common structural factors with the original molecules, hopefully making them good starting points for combinatorial optimization within the chemical space (Gao et al., 2021; Swanson et al., 2023).
@@ 4.4 Projecting Molecules to Explore Local Chemical Space for Hit Expansion
Our model’s capabilities also extend to hit expansion in drug discovery, offering a strategic approach to identifying structurally similar, synthesizable analogs of hit molecules (Keserű & Makara, 2006; Levin et al., 2023).
We showcase our model’s hit expansion capability by designing inhibitors for c-Jun N-terminal Kinases-3 (JNK3), following the setting of Levin et al.. Based on the oracle function provided by Li et al., our goal is to identify high-scoring, synthesizable molecules similar to a specified hit molecule with a JNK3 score of 0.68. We generated 500 analogs for the hit, resulting in a diverse array of analogs with different JNK3 scores and structural similarities, as depicted in Figure 5. Remarkably, 21 analogs outperformed the original hit in JNK3 score while maintaining an average Tanimoto similarity of 0.67. Additionally, these analogs present distinct synthetic pathways, different from traditional synthesis planning and building block exchange strategies (Levin et al., 2023), underlining our model’s capability to cover broader chemical space (Figure 8 in the appendix).
@@ A.1 Training Data
  Global: ${\mathcal{H}}:R\mapsto\{M\}$, a hash map that stores eligible molecules for each reaction.
@@ A.2 Parameters
The molecular graph encoder consists of 8 graph transformer layers, each of which has 8 attention heads, and the dimension of the input and output features is 512. The postfix notation decoder is a stack of 8 transformer decoder layers. Each has 8 attention heads, and the dimension of the input and output features is also 512. In the nearest-neighbor search, each building block molecule is indexed by the Morgan fingerprint (Morgan, 1965) of length 256 and radius 2. We use the AdamW optimizer (Loshchilov & Hutter, 2017) to train the network with a learning rate of 3e-4 and a batch size of 256 for 500,000 iterations. During the construction of synthesis for training, we limit the maximum number of reactions to 5 and the maximum number of atoms to 80. At inference time, we set the maximum sequence length to 16. If the end token is not generated within the limit, the process will be considered unsuccessful.
@@ B.1 Stricter Test Split for Bottom-Up Synthesis Planning
Method SuccessRecons.Sim.(Morgan) Sim.(Scaffold) Sim.(Gobbi) SynNet 84.0% 10.0% 0.4552 0.6012 0.3348 Proposed 95.8% 17.0% 0.5921 0.7239 0.6110
@@ B.3 Baseline for Hit Expansion
In hit expansion, the objective is to identify similar molecules within the local chemical space. Given that random molecular edits may not guarantee synthesizability, we adopt querying a chemical database as the baseline method. Specifically, we search PubChem (Kim et al., 2023) using the hit molecule and retrieve the most similar 500 molecules based on fingerprint similarity. Only 7 out of the retrieved molecules surpass the hit in JNK3 score, with an average similarity score of 0.67. This result underscores the efficiency of our model in identifying relevant synthesizable compounds within the chemical space.

### v2

@@ 3.5 Experiment Setup
We use the SynNet reaction template set (gao2021synnet) curated by the authors from two publicly available template sets (hartenfeller2012dogs; button2019automated). The template set contains 91 reaction templates encoded in SMARTS string format (DaylightTheory2023), of which 13 are uni-molecular reactions and 78 are bi-molecular reactions.
@@
We use the building blocks in the Enamine US Stock catalog retrieved in October 2023 (enamine). For building blocks that include more than one molecule (e.g., salts, hydrates), we keep the largest molecule and drop the remaining ones. Building blocks that fail the RDKit sanitization check or do not match any reaction template are removed. Duplicate building blocks are also removed. The pre-processing procedure leads to the final building block set containing 211,220 molecules.
@@
To evaluate the similarity between the input molecule and the output molecule, we use the Tanimoto similarity score on three different fingerprints: (1) Morgan fingerprint of length 4096 and radius 2 (morgan1965generation), (2) Morgan fingerprint of Murcko scaffold, and (3) Gobbi pharmacophore fingerprint (gobbi1998genetic). The three similarity scores indicate chemical similarities in three different aspects: overall structure, scaffold structure, and pharmacophore property, respectively, and are all normalized to $[0,1]$.
@@ 4.1 Generalization in Bottom-Up Synthesis Planning
We first validate our model’s ability in bottom-up synthesis planning, particularly in proposing synthetic pathways for novel molecular structures. To assess this capability with unseen molecules, we employed the following testing approach. First, we clustered building blocks into 128 groups using K-means algorithm based on their fingerprints, reserving one cluster exclusively for testing while using the remaining 127 for training. This approach generated a test set of 1,000 molecules from the reserved building block to evaluate our model’s generalization. During testing, we allow the model to access the reserved building blocks. In addition, we include another challenging benchmark: 1,000 molecules from the ChEMBL database (Release 33)(gaulton2012chembl), reported to be predominantly “unreachable” in previous work (gao2021synnet). This not only tested our model against complex, unseen molecular structures but also set a strenuous standard for assessing its generalizability.
We evaluate the model using the following metrics: (1) success rate: the percentage of valid postfix notations; (2) reconstruction rate: the percentage of proposed synthetic pathways that lead to products identical to input molecules; (3) average similarities. The similarity score is set to 0 for failed cases. We compare our model with the baseline model SynNet (gao2021synnet).
@@ 4.2 Projecting Molecules Generated by Structure-Based Drug Design Models
Generative models for structure-based drug design (SBDD) build molecules to fit a protein pocket structure (luo20213sbdd). SBDD models often produce chemically invalid structures, because they operate in the 3D space with a few chemical constraints (harris2023benchmarksbdd). As a remedy, we apply our model to molecules proposed by the SBDD models to find valid and synthesizable analogs.
We use Pocket2Mol (peng2022pocket2mol) to generate molecules for targets in the LIT-PCBA dataset (tran2020litpcba). The LIT-PCBA dataset contains 15 protein targets, each of which has at least one X-ray structure that can be used to run Pocket2Mol to generate molecules. For each target, we select 300 generated molecules based on the sum of QED and SA scores. Then, we apply our model to generate 5 analogs for each molecule and select best one according to Vina scores (eberhardt2021autodock). The estimated binding energy in kcal/mol shows slight improvement for the majority of targets (Table 3), suggesting that the top-1 synthesizable analogs of each molecule could maintain binding. Figure 4 presents two examples in which the synthesizable analogs improve predicted binding. Figure 6 and 7 in the appendix present more examples.
@@ 4.3 Projecting Molecules Generated by Goal-Directed Generative Models
Goal-directed generative models are pivotal in molecular design, yet their practicality is often hampered by the synthesizability of the produced molecules. This issue was highlighted by gao2020synthesizability, who, utilizing three algorithms on the GuacaMol benchmark suite’s multi-property objectives (brown2019guacamol), revealed that around 70% of the molecules generated were considered unsynthesizable by the ASKCOS retrosynthesis analysis tool (coley2019askcos).
@@
In conclusion, the projection generally leads to synthesizable analogs that are sub-optimal in terms of the oracle property scores. These sub-optimal analogs share some common structural factors with the original molecules, hopefully making them good starting points for combinatorial optimization within the chemical space (gao2021synnet; swanson2023generative).
@@ 4.4 Projecting Molecules to Explore Local Chemical Space for Hit Expansion
Our model’s capabilities also extend to hit expansion in drug discovery, offering a strategic approach to identifying structurally similar, synthesizable analogs of hit molecules (keserHu2006hit; levin2023hitexpand).
We showcase our model’s hit expansion capability by designing inhibitors for c-Jun N-terminal Kinases-3 (JNK3), following the setting of levin2023hitexpand. Based on the oracle function provided by li2018multi, our goal is to identify high-scoring, synthesizable molecules similar to a specified hit molecule with a JNK3 score of 0.68. We generated 500 analogs for the hit, resulting in a diverse array of analogs with different JNK3 scores and structural similarities, as depicted in Figure 5. Remarkably, 21 analogs outperformed the original hit in JNK3 score while maintaining an average Tanimoto similarity of 0.67. Additionally, these analogs present distinct synthetic pathways, different from traditional synthesis planning and building block exchange strategies (levin2023hitexpand), underlining our model’s capability to cover broader chemical space (Figure 8 in the appendix).
@@ A.1 Training Data
Global: ${\mathcal{H}}:R\mapsto\{M\}$, a function that stores available building blocks for each reaction.
@@ A.2 Parameters
The molecular graph encoder consists of 8 graph transformer layers, each of which has 8 attention heads, and the dimension of the input and output features is 512. The postfix notation decoder is a stack of 8 transformer decoder layers. Each has 8 attention heads, and the dimension of the input and output features is also 512. In the nearest-neighbor search, each building block molecule is indexed by the Morgan fingerprint (morgan1965generation) of length 256 and radius 2. We use the AdamW optimizer (loshchilov2017adamw) to train the network with a learning rate of 3e-4 and a batch size of 256 for 500,000 iterations. During the construction of synthesis for training, we limit the maximum number of reactions to 5 and the maximum number of atoms to 80. At inference time, we set the maximum sequence length to 16. If the end token is not generated within the limit, the process will be considered unsuccessful.
@@ B.1 Stricter Test Split for Bottom-Up Synthesis Planning
Method Success% Recons.% Sim.(Morgan) Sim.(Scaffold) Sim.(Gobbi) SynNet 84.0% 10.0% 0.4552 0.6012 0.3348 ChemProjector 95.8% 17.0% 0.5921 0.7239 0.6110
@@ B.3 Baseline for Hit Expansion
In hit expansion, the objective is to identify similar molecules within the local chemical space. Given that random molecular edits may not guarantee synthesizability, we adopt querying a chemical database as the baseline method. Specifically, we search PubChem (kim2023pubchem) using the hit molecule and retrieve the most similar 500 molecules based on fingerprint similarity. Only 7 out of the retrieved molecules surpass the hit in JNK3 score, with an average similarity score of 0.67. This result underscores the efficiency of our model in identifying relevant synthesizable compounds within the chemical space.